# Language Agents for Hypothesis-driven Clinical Decision Making with Reinforcement Learning

### v1

@@ Abstract
Clinical decision-making is a dynamic, interactive, and cyclic process where doctors have to repeatedly decide on which clinical action to perform and consider newly uncovered information for diagnosis and treatment. Large Language Models (LLMs) have the potential to support clinicians in this process, however, most applications of LLMs in clinical decision support suffer from one of two limitations: Either they assume the unrealistic scenario of immediate availability of all patient information and do not model the interactive and iterative investigation process, or they restrict themselves to the limited "out-of-the-box" capabilities of large pre-trained models without performing task-specific training. In contrast to this, we propose to model clinical decision-making for diagnosis with a hypothesis-driven uncertainty-aware language agent, LA-CDM, that converges towards a diagnosis via repeatedly requesting and interpreting relevant tests. Using a hybrid training paradigm combining supervised and reinforcement learning, we train LA-CDM with three objectives targeting critical aspects of clinical decision-making: accurate hypothesis generation, hypothesis uncertainty estimation, and efficient decision-making. We evaluate our methodology on MIMIC-CDM, a real-world dataset covering four abdominal diseases containing various clinical tests and show the benefit of explicitly training clinical decision-making for increasing diagnostic performance and efficiency.
@@ 1 Introduction
In modern healthcare systems, accurate and timely diagnosis stands as a critical component in patient management and treatment Savioli et al. (2022). To form a diagnosis, clinicians often engage in clinical decision-making through a dynamic, iterative process, called differential diagnosis. This process involves forming hypotheses about the patient and testing those hypotheses through requesting and interpreting information from relevant available diagnostic tests Sox et al. (2024). At the beginning multiple diagnoses may be possible and there is still high uncertainty. Through diagnostic testing the uncertainty is minimized and the space of possible diseases reduced until a sufficient confidence is reached, a diagnosis can be given, and treatment can begin Sox et al. (2024). It is especially important in complex, high-stakes environments such as emergency departments, where often little is known about a patient at admission and fast and accurate diagnosis is paramount Savioli et al. (2022).
@@
Large Language Models (LLMs), with their ability to synthesize complex textual information, seem well-suited for aiding clinicians in this task, as most medical information is textual or can be represented in text, e.g., clinical notes, imaging reports or numerical laboratory results. This enables great flexibility and variety in medical modalities. In the medical domain LLMs have already shown great success in passing medical license exams Singhal et al. (2023); Gilson et al. (2023) and diagnosing case challenges Buckley et al. (2023). The potential to revolutionize healthcare through accurate diagnostic capabilities is enormous, however, existing approaches often either (1) assume immediate availability of all patient data Buckley et al. (2023); McDuff et al. (2023); Chen et al. (2025), which is rarely the case in practice, or (2) rely on the often limited “out-of-the-box” behavior of pre-trained LLMs without any task-specific fine-tuning to the complexities of diagnostic decision-making Hager et al. (2024); Li et al. (2024). This mismatch between research and real-world clinical decision-making limits the applicability of LLMs to the clinical setting.
In this paper, we address the above limitations by modeling and training Language Agents for Clinical Decision Making (LA-CDM), tasked with iteratively reducing hypothesis uncertainty through repeated diagnostic testing. Inspired by cognitive research on human clinical decision-making Sox et al. (2024), we design a two-agent system replicating the two main cognitive tasks of clinicians involved in clinical decision-making. It consists of one LLM agent, the hypothesis agent, forming the most likely diagnosis hypothesis based on all available patient information and estimating its confidence in that hypothesis, and another agent, the decision agent, that evaluates the patient information and the hypothesis agent’s output to either provide a diagnosis or request an additional diagnostic test (Figure 1). To train this system, we propose a novel training strategy with three distinct objectives that target the core principles of successful clinical decision-making Sox et al. (2024):
@@
    Efficient decision-making: Using reinforcement learning, the decision agent is trained to select the most informative next test and reach a diagnosis when sufficiently confident. The model gets rewarded for a final correct diagnosis, reinforcing the testing pathway that led to that diagnosis.
Analogously, to doctors graduating from medical school, LLMs have a strong medical knowledge foundation, but are not trained on performing clinical decision-making. Clinicians learn this skill through years of experience, pointing towards experience-based reinforcement learning as a prime paradigm for teaching clinical decision-making to LLMs. Further, since optimal testing pathways are not known, supervised learning of clinical decision-making is infeasible. The interplay of the three objectives results in the model learning which tests to request in order to increase the hypothesis confidence leading to an accurate diagnosis. This guides the model to request those tests that are most informative in a given situation. To the best of our knowledge, we propose the first method for explicitly training LLMs for clinical decision-making.
We evaluate our method on MIMIC-CDM Hager et al. (2024), a real-world dataset focused on four abdominal diseases that mirrors clinical workflows by modeling differential diagnosis through sequential test requests. Its standardized test naming and inclusion of lab results, notes, and imaging reports make it uniquely suited for training and evaluating LLM-based diagnostic reasoning. Our experiments demonstrate the benefit of explicitly training clinical decision-making. Notably, training reduces the required diagnostic tests, which has practical implications in reducing healthcare costs, diagnosis time, and patient discomfort Savioli et al. (2022). We show the benefit of our hypothesis-driven approach to clinical decision-making and demonstrate that the model adapts its testing procedure to the patient at hand, placing this work as a step towards patient-specific personalized differential diagnosis.
@@ 2 Related Work
ED-Copilot Sun et al. (2024) employs a Language Model for encoding serialized patient laboratory values which is trained end-to-end with two MLPs, one predicting test requests, and one predicting the task of outcome severity. They train their method in two stages. First, they use supervised learning to teach the model to predict all tests in a pre-defined order. In a second training stage, they use reinforcement learning to finetune the model to reduce time cost of the testing regiment. They also report great reduction in testing time, while remaining similar in performance to baselines making use of all available information. While this method uses a language model as its encoding backbone, this language model is not directly requesting tests. Same as the previous method, they further only consider tabular laboratory values and therefore do not make use of valuable information in textual form.
Reflexion Shinn et al. (2023) addresses general LLM decision-making by introducing a zero-shot reinforcement learning approach. Rather than fine-tuning the LLM with traditional reinforcement learning, the method allows the model to iteratively attempt the same task. After each trial, a limited number of previous attempts and a numerical reward are added to the input context, guiding the next generation. While this is effective at general decision-making, this approach cannot be applied in the clinical context, where multiple diagnosis trials with the same patient and intermediate diagnosis correctness feedback are infeasible.
LLMs for Clinical Decision Making   LLMs have so far only been used as zero-shot methods for clinical decision support. Hager et al. (2024) place large "out-of-the-box" language models in an evaluation framework where they are tasked with interactively requesting diagnostic tests and diagnose patients. They show severe limitations of LLMs for clinical decision-making and report worse diagnostic performance than clinicians. Vaid et al. (2024) approach clinical decision-making with a tool-using LLM. Through zero-shot prompting, they provide the LLM with a number of available tools, e.g., a symptom tool for retrieving the symptoms, or a imaging study tool for getting any imaging reports on the patient. They evaluate various proprietary LLMs, with GPT-4 Achiam et al. (2023) showing the best performance. Liu et al. (2024) model LLM clinical decision-making as a multi-agent setting, where a doctor agent communicates with a patient agent who can detail his symptoms and a technician agent who can perform laboratory or imaging results. They compare different prompting techniques, like chain-of-thought Wei et al. (2022) or one-shot prompting and show that GPT-4o achieves the best performance. All these methods do not attempt to improve model performance by explicitly training clinical decision-making with LLMs.
@@ 3.1 Reinforcement Learning on Large Language Models
Reinforcement learning has emerged as a powerful framework for fine-tuning LLMs by aligning their behavior with desired objectives, especially when no clear ground truth for this behaviour exists. Most notably, Reinforcement Learning with Human Feedback (RLHF) Ouyang et al. (2022) was used to align LLM generations with human preferences. LA-CDM utilizes a direct reinforcement learning method based on Proximal Policy Optimization (PPO) Schulman et al. (2017) without requiring human feedback.
@@ 3.2 Confidence Calibration of Large Language Models
While LLMs have shown impressive capabilities in many language-related tasks, hallucinations and confidently-presented wrong answers are a common and well-known problem Kadavath et al. (2022). A well-calibrated model is able to express confidence that aligns with the epistemic probability of correctness. This means that of all the answers which are presented with a confidence of $0\leq p\leq 1$, the fraction of correct answers is $p$.
@@
where $y_{pred}$ is the predicted answer, $10<c<$ 1 is the (scaled and clipped) confidence prediction, and $J(\cdot)$ is a binary function evaluating the correctness of $y_{pred}$. The reward is then scaled to be between $-1$ and $1$. They train the model using Proximal Policy Optimization (PPO) Schulman et al. (2017). This training approach removes the need for an artificially constructed ground truth confidence dataset, as done by other confidence calibration methods Azaria and Mitchell (2023); Kadavath et al. (2022), and instead only requires a measure of answer correctness. The authors prove that an optimal policy under their reward design produces perfectly calibrated confidence expressions.
@@ 4.1 Modeling Clinical Decision Making
Clinical Decision Making Environment   In order to train our model in decision-making it has to operate in a reinforcement learning environment to explore testing strategies and receive reward signals. The model learns decision-making through interacting with this environment and diagnosing patients. Let each patient be described by a number of $n$ test results $[t_{i}]^{n}_{i=1}$ as textual records of clinical notes, imaging reports and laboratory panels. Since patient information is iteratively uncovered through the model’s test requests $r_{j}$, we define the currently observed patient state at time-step $j$ as the set of all observed tests and write it as $p_{j}$. The correct diagnosis for the patient is denoted by $y_{true}$.
@@
The model generates this output in a format of "Hypothesis: $h_{j}$, Confidence: $c_{j}$". The agent reports numerical confidence estimations on a scale of 00 to $10$, where $10$ means absolute certainty of the correctness of the hypothesis, and 00 means absolute certainty that it is incorrect.
Decision Agent   The decision agent $\mathcal{D}$ is the actor advancing the environment. It decides on which action to take at each time-step. Through its system prompt it is provided with its task, a list of tests present in the dataset, and the possible patient diagnoses. Provided with the currently observed patient state $p_{j}$ and the hypothesis agent’s hypothesis $h_{j}$ and confidence $c_{j}$, it produces a decision on whether to request another diagnostic test $r_{j}$ and move on to the next time-step $j+1$ or whether to commit on a specific diagnosis $y_{pred}$ for the patient and end the episode. Formally, it produces a mapping
$\mathcal{D}:\{p_{j},h_{j},c_{j}\}\xrightarrow{}\begin{cases}r_{j}&\textit{if afurther test is requested}\\ y_{pred}&\textit{if a diagnosis is given}\end{cases}$
Specifically, we employ the ReAct prompting technique Yao et al. (2022), to prime the model to first produce a reasoning trace, following chain-of-thought principles Wei et al. (2022), and then provide action and action input (in our case, the specific test or diagnosis) in a structured format.
@@ 4.2 Training Clinical Decision Making
Training Uncertainty-Awareness   Since the available patient information is often limited, especially at early stages of the diagnostic process, and the available data does not always clearly point at a specific diagnosis, uncertainty is inherent to clinical decision-making. An accurate intrinsic estimation of that uncertainty by the model can give it an improved basis for decisions on when to stop the diagnostic process and produce a diagnosis. In this work, we train confidence calibration following a method proposed by Stangel et al. (2025) as previously introduced in Section 3.2. We define our correctness measure $J(h_{j})$ as equality between the predicted hypothesis $h_{j}$ and the ground truth diagnosis $y_{true}$. Since "out-of-the-box" pre-trained LLMs tend to only predict high confidences, the model rarely explores predicting low confidences, significantly hindering the reinforcement learning training. We take inspiration from stochastic reinforcement learning exploration techniques and, with probability $p_{explore}$, replace the predicted confidence with a randomly chosen different confidence. The exploration probability is scheduled to decrease during training.
@@
Training Clinical Action Selection   At the core of training clinical decision-making lies the training of clinical action selection. During interaction with the clinical decision-making environment the model can freely choose to iteratively request any number of tests in any order. Given the vast number of possible diagnostic pathways, defining an optimal test sequence for each patient is infeasible, we therefore do not have a ground truth on which tests to perform. To still be able to train clinical decision-making, we propose to use reinforcement learning, where the model can learn through trial-and-error which tests are useful in which situations. Equal to the confidence calibration training, we employ the PPO algorithm Schulman et al. (2017) for reinforcement learning training. Through interaction with the clinical decision-making environment the model requests different tests until it decides on a diagnosis. We design our reward function to present the model with a fixed positive reward $r_{pos}$ if the final diagnosis at the end of the diagnosis episode was correct, or a fixed negative reward $r_{neg}$ if it is wrong. Additionally, we punish the model with reward $r_{invalid}$ if the model violates the specified format. Our reward function is thus:
@@
$R(y_{pred})=\begin{cases}r_{pos}&\textit{if $y_{pred}=y_{true}$}\\ r_{neg}&\textit{if $y_{pred}\neq y_{true}$}\\ r_{invalid}&\textit{for out-of-format generations}\end{cases}$
@@
The interaction of these three objectives enables the model to learn which tests to request to enhance hypothesis confidence, ultimately leading to a more accurate diagnosis. This drives the model to prioritize tests that provide the most informative insights in a given situation.

### v2

@@ Abstract
Clinical decision-making is a dynamic, interactive, and cyclic process where doctors have to repeatedly decide on which clinical action to perform and consider newly uncovered information for diagnosis and treatment. Large Language Models (LLMs) have the potential to support clinicians in this process, however, most applications of LLMs in clinical decision support suffer from one of two limitations: Either they assume the unrealistic scenario of immediate availability of all patient information and do not model the interactive and iterative investigation process, or they restrict themselves to the limited ”out-of-the-box” capabilities of large pre-trained models without performing task-specific training. In contrast to this, we propose to model clinical decision-making for diagnosis with a hypothesis-driven uncertainty-aware language agent, LA-CDM, that converges towards a diagnosis via repeatedly requesting and interpreting relevant tests. Using a hybrid training paradigm combining supervised and reinforcement learning, we train LA-CDM with three objectives targeting critical aspects of clinical decision-making: accurate hypothesis generation, hypothesis uncertainty estimation, and efficient decision-making. We evaluate our methodology on MIMIC-CDM, a real-world dataset covering four abdominal diseases containing various clinical tests and show the benefit of explicitly training clinical decision-making for increasing diagnostic performance and efficiency. Our code is available at https://github.com/dharouni/LA-CDM.
@@ 1 Introduction
In modern healthcare systems, accurate and efficient diagnosis stands as a critical component in patient management and treatment (Savioli et al., 2022). To form a diagnosis, clinicians often engage in clinical decision-making through a dynamic, iterative process, called differential diagnosis. This process involves forming hypotheses about the patient and testing those hypotheses through requesting and interpreting information from relevant available diagnostic tests (Sox et al., 2024). At the beginning multiple diagnoses may be possible and there is still high uncertainty. Through diagnostic testing the uncertainty is minimized and the space of possible diseases reduced until a sufficient confidence is reached, a diagnosis can be given, and treatment can begin (Sox et al., 2024). It is especially important in complex, high-stakes environments such as emergency departments, where often little is known about a patient at admission and fast and accurate diagnosis is paramount (Savioli et al., 2022).
@@
Large Language Models (LLMs), with their ability to synthesize complex textual information, seem well-suited for aiding clinicians in this task, as most medical information is textual or can be represented in text, e.g., clinical notes, imaging reports or numerical laboratory results. This enables great flexibility and variety in medical modalities. In the medical domain LLMs have already shown great success in passing medical license exams (Singhal et al., 2023; Gilson et al., 2023) and diagnosing case challenges (Buckley et al., 2023). The potential to revolutionize healthcare through accurate diagnostic capabilities is enormous, however, existing approaches often either (1) assume immediate availability of all patient data (Buckley et al., 2023; McDuff et al., 2023; Chen et al., 2025), which is rarely the case in practice, or (2) rely on the often limited “out-of-the-box” behavior of pre-trained LLMs without any task-specific fine-tuning to the complexities of diagnostic decision-making (Hager et al., 2024; Li et al., 2024; Nori et al., 2025). This mismatch between research and real-world clinical decision-making limits the applicability of LLMs to the clinical setting.
In this paper, we address the above limitations by modeling and training Language Agents for Clinical Decision Making (LA-CDM), tasked with iteratively reducing hypothesis uncertainty through repeated diagnostic testing. Inspired by cognitive research on human clinical decision-making (Sox et al., 2024), we design a two-agent system replicating the two main cognitive tasks of clinicians involved in clinical decision-making. It consists of one LLM agent, the hypothesis agent, forming the most likely diagnosis hypothesis based on all available patient information and estimating its confidence in that hypothesis, and another agent, the decision agent, that evaluates the patient information and the hypothesis agent’s output to either provide a diagnosis or request an additional diagnostic test (Figure 1). To train this system, we propose a novel training strategy with three distinct objectives that target the core principles of successful clinical decision-making (Sox et al., 2024):
@@
    Efficient decision-making: Using reinforcement learning, the decision agent is trained to select the most informative next test and reach a diagnosis when sufficiently confident, taking test costs into account. The model gets rewarded for a final correct diagnosis, reinforcing the testing pathway that led to that diagnosis.
Analogously to doctors graduating from medical school, LLMs have a strong medical knowledge foundation, but are not trained on performing clinical decision-making. Clinicians learn this skill through years of experience, pointing towards experience-based reinforcement learning as a prime paradigm for teaching clinical decision-making to LLMs. Further, since optimal testing pathways are not known, supervised learning of clinical decision-making is infeasible. The interplay of the three objectives results in the model learning which tests to request in order to increase the hypothesis confidence leading to an accurate diagnosis at minimal testing cost. This guides the model to request those tests that are most informative in a given situation. To the best of our knowledge, we propose the first method for explicitly training LLMs for clinical decision-making.
We evaluate our method on MIMIC-CDM (Hager et al., 2024), a real-world dataset focused on four abdominal diseases that mirrors clinical workflows by modeling differential diagnosis through sequential test requests. Its standardized test naming and inclusion of lab results, notes, and imaging reports make it uniquely suited for training and evaluating LLM-based diagnostic reasoning. Our experiments demonstrate the benefit of explicitly training clinical decision-making. Notably, training reduces the diagnostic cost, which has practical implications in reducing healthcare costs, diagnosis time, and patient discomfort (Savioli et al., 2022). We show the benefit of our hypothesis-driven approach to clinical decision-making and demonstrate that the model adapts its testing procedure to the patient at hand, placing this work as a step towards patient-specific personalized differential diagnosis.
@@ 2 Related Work
ED-Copilot (Sun et al., 2024) employs a Language Model for encoding serialized patient laboratory values which is trained end-to-end with two MLPs, one predicting test requests, and one predicting the task of outcome severity. They train their method in two stages. First, they use supervised learning to teach the model to predict all tests in a pre-defined order. In a second training stage, they use reinforcement learning to finetune the model to reduce time cost of the testing regiment. They also report great reduction in testing time, while remaining similar in performance to baselines making use of all available information. While this method uses a language model as its encoding backbone, this language model is not directly requesting tests. Same as the previous method, they further only consider tabular laboratory values and therefore do not make use of valuable information in textual form.
Reflexion (Shinn et al., 2023) addresses general LLM decision-making by introducing a zero-shot reinforcement learning approach. Rather than fine-tuning the LLM with traditional reinforcement learning, the method allows the model to iteratively attempt the same task. After each trial, a limited number of previous attempts and a numerical reward are added to the input context, guiding the next generation. While this is effective at general decision-making, this approach cannot be applied in the clinical context, where multiple diagnosis trials with the same patient and intermediate diagnosis correctness feedback are infeasible.
LLMs for Clinical Decision Making   LLMs have so far only been used as zero-shot methods for clinical decision support. Hager et al. (2024) place large ”out-of-the-box” language models in an evaluation framework where they are tasked with interactively requesting diagnostic tests and diagnose patients. They show severe limitations of LLMs for clinical decision-making and report worse diagnostic performance than clinicians. Vaid et al. (2024) approach clinical decision-making with a tool-using LLM. Through zero-shot prompting, they provide the LLM with a number of available tools, e.g., a symptom tool for retrieving the symptoms, or a imaging study tool for getting any imaging reports on the patient. They evaluate various proprietary LLMs, with GPT-4 (Achiam et al., 2023) showing the best performance. Liu et al. (2024) model LLM clinical decision-making as a multi-agent setting, where a doctor agent communicates with a patient agent who can detail his symptoms and a technician agent who can perform laboratory or imaging results. They compare different prompting techniques, like chain-of-thought (Wei et al., 2022) or one-shot prompting and show that GPT-4o achieves the best performance. All these methods do not attempt to improve model performance by explicitly training clinical decision-making with LLMs.
@@ 3.1 Reinforcement Learning on Large Language Models
Reinforcement learning has emerged as a powerful framework for fine-tuning LLMs by aligning their behavior with desired objectives, especially when no clear ground truth for this behaviour exists. Most notably, Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) was used to align LLM generations with human preferences. LA-CDM utilizes a direct reinforcement learning method based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024) without requiring human feedback.
@@ 3.2 Confidence Calibration of Large Language Models
While LLMs have shown impressive capabilities in many language-related tasks, hallucinations and confidently-presented wrong answers are a common and well-known problem (Kadavath et al., 2022). A well-calibrated model is able to express confidence that aligns with the epistemic probability of correctness. This means that of all the answers which are presented with a confidence of $0\leq p\leq 1$, the fraction of correct answers is $p$.
@@
where $y_{pred}$ is the predicted answer, $0<c<1$ is the (scaled and clipped) confidence prediction, and $J(\cdot)$ is a binary function evaluating the correctness of $y_{pred}$. The reward is then scaled to be between 1-1 and $1$. They train the model using Proximal Policy Optimization (PPO) (Schulman et al., 2017). This training approach removes the need for an artificially constructed ground truth confidence dataset, as done by other confidence calibration methods (Azaria & Mitchell, 2023; Kadavath et al., 2022), and instead only requires a measure of answer correctness. The authors prove that an optimal policy under their reward design produces perfectly calibrated confidence expressions.
@@ 4.1 Modeling Clinical Decision Making
Clinical Decision Making Environment   In order to train our model in decision-making it has to operate in a reinforcement learning environment to explore testing strategies and receive reward signals. The model learns decision-making through interacting with this environment and diagnosing patients. Let each patient be described by a number of $n$ test results $t_{i}]^{n}_{i=1}$ as textual records of clinical notes, imaging reports and laboratory panels. Since patient information is iteratively uncovered through the model’s test requests $t_{j}$, we define the currently observed patient state at time-step $j$ as the set of all observed tests and write it as $p_{j}$. The correct diagnosis for the patient is denoted by $y_{true}$.
@@
The model generates this output in a format of ”Hypothesis: $h_{j}$, Confidence: $c_{j}$”. The agent reports numerical confidence estimations on a scale of 0 to $10$, where $10$ means absolute certainty of the correctness of the hypothesis, and 0 means absolute certainty that it is incorrect.
Decision Agent   The decision agent $\mathcal{D}$ is the actor advancing the environment. It decides on which action to take at each time-step. Through its system prompt it is provided with its task, a list of tests present in the dataset, and the possible patient diagnoses. Provided with the currently observed patient state $p_{j}$ and the hypothesis agent’s hypothesis $h_{j}$ and confidence $c_{j}$, it produces a decision on whether to request another diagnostic test $t_{j}$ and move on to the next time-step $j+1$ or whether to commit on a specific diagnosis $y_{pred}$ for the patient and end the episode. Formally, it produces a mapping
[…]
Specifically, we employ the ReAct prompting technique (Yao et al., 2022), to prime the model to first produce a reasoning trace, following chain-of-thought principles (Wei et al., 2022), and then provide action and action input (in our case, the specific test or diagnosis) in a structured format.
@@ 4.2 Training Clinical Decision Making
Training Uncertainty-Awareness   Since the available patient information is often limited, especially at early stages of the diagnostic process, and the available data does not always clearly point at a specific diagnosis, uncertainty is inherent to clinical decision-making. An accurate intrinsic estimation of that uncertainty by the model can give it an improved basis for decisions on when to stop the diagnostic process and produce a diagnosis. In this work, we train confidence calibration following a method proposed by Stangel et al. (2025) as previously introduced in Section 3.2, however instead of PPO, we use the GRPO algorithm (Shao et al., 2024). We define our correctness measure $J(h_{j})$ as equality between the predicted hypothesis $h_{j}$ and the ground truth diagnosis $y_{true}$.
@@
Training Efficient Clinical Action Selection   At the core of training clinical decision-making lies the training of clinical action selection. During interaction with the clinical decision-making environment the model can freely choose to iteratively request any number of tests in any order. Given the vast number of possible diagnostic pathways, defining an optimal test sequence for each patient is infeasible, we therefore do not have a ground truth on which tests to perform. To still be able to train clinical decision-making, we propose to use reinforcement learning, where the model can learn through trial-and-error which tests are useful in which situations. Equal to the confidence calibration training, we employ the GRPO algorithm (Shao et al., 2024) for reinforcement learning training. Through interaction with the clinical decision-making environment the model freely requests different tests until it decides on a diagnosis. The requested tests are provided to the model by the environment whenever available. If not available the model is notified and asked to request a different test. We design our diagnosis reward function $R_{diag}$ to present the model with a fixed positive reward $r_{pos}$ if the final diagnosis at the end of the diagnosis episode was correct, or a fixed negative reward $r_{neg}$ if it is wrong. Additionally, we punish the model with reward $r_{invalid}$ if the model violates the specified format. Our reward function is thus:
@@
$R_{diag}(y_{pred})=\begin{cases}r_{pos}&\textit{if $y_{pred}=y_{true}$}\\ r_{neg}&\textit{if $y_{pred}\neq y_{true}$}\\ r_{invalid}&\textit{for out-of-format generations}\end{cases}$
Not just diagnosis accuracy but also diagnostic efficiency is an important factor in clinical decision making. Different tests have different costs for the healthcare system, e.g., a CT scan is much more expensive, than a simple blood value. The model should therefore only request expensive tests, when necessary for diagnosis. To incentivize efficient behaviour, we add an additional reward function $R_{cost}$ that punishes test usage relative to their cost:
$R_{cost}(T)=-\sum_{t_{j}\in T}c(t_{j}),$
where $T$ is the list of all performed tests and $c(t_{j})$ is the cost of test $t_{j}$.
@@
The interaction of these three objectives enables the model to learn which tests to request to enhance hypothesis confidence, ultimately leading to a more accurate diagnosis while also taking test cost into account. This drives the model to prioritize tests that provide the most informative insights in a given situation.